OpenStudy (anonymous):

10^4√6+4^3√6-9^3√6-3^4√6 Answers: 2√6 7^4√6-5^3√6 7^4√12-5^3√6 2√12

OpenStudy (saifoo.khan):

\[\large 10^4 \sqrt(6)+4^3 \sqrt(6)-9^3 \sqrt(6)-3^4 \sqrt(6) \] like this?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

is it 10^4 or fourthroot?

OpenStudy (anonymous):

fourthroot

OpenStudy (saifoo.khan):

please use Equation editor and write it.

OpenStudy (anonymous):

\[10\sqrt[4]{6}+4\sqrt[3]{6}-9\sqrt[3]{6}-3\sqrt[4]{6}\] \[7\sqrt[4]{6}-5\sqrt[3]{6}\]

OpenStudy (anonymous):

thank you!

OpenStudy (saifoo.khan):

righto!